the sum of three consecutiveintegers is 48. what is the set of three consecutive integers that fitsthis situation

Answer by mananth(16949)

let the integers be x, x+1, x+2
..
x+x+1+x+2=48
3x+3=48
add -3
3x=48-3
3x=45
/3
x=15
..
15,16,17 are the integers
..
m.ananth@hotmail.ca